# Video LLMs for Temporal Reasoning in Long Videos

### v4

@@ A.8 Impacts of Number of Short-Term Clips
Table 11: Effects of the number of short-term clips. Best results are in bold, while second best ones are underlined.
A.9 General Video Understanding Results
Method Data CI DO CU TU C
Video-LLaMA [57] 10M 1.96 2.18 2.16 1.82 1.79
Video-ChatGPT [29] 100k 2.50 2.57 2.69 2.16 2.20
Valley [28] 234k 2.43 2.13 2.86 2.04 2.45
BT-Adapter [25] 10M 2.68 2.69 3.27 2.34 2.46
LongVLM [50] 100k 2.76 2.86 3.34 2.39 3.11
TemporalVLM 100k 2.88 2.91 3.45 2.50 3.16
Following LongVLM [50], we evaluate the performance of our TemporalVLM model on the general video understanding benchmark provided by Video-ChatGPT [29] in Tab. 12. The evaluation metrics include Correctness Information (CI), Detail Orientation (DO), Contextual Understanding (CU), Temporal Understanding (TU), and Consistency (C). It is evident from Tab. 12 that our TemporalVLM model achieves the best performance across all metrics, outperforming all competing methods, including the original LongVLM model [50].
A.10 Qualitative Results
In addition to Figs. 4 and 5 in the main paper, we provide additional qualitative results in this section. In particular, Figs. 7, 8, and 9 present qualitative results by TemporalVLM in the supervised setting for dense video captioning, video highlight detection, and temporal video grounding respectively. More importantly, qualitative results demonstrating the generalization abilities of TemporalVLM in the zero-shot setting are shown in Fig. 10. Overall, our TemporalVLM model demonstrates promising performance in a variety of temporal reasoning and fine-grained understaning tasks.
@@ A.10 Qualitative Results
A.11 Discussions
A.12 Limitations
Despite promising performance in several temporal reasoning tasks, including dense video captioning, temporal video grounding, video highlight detection, and temporal action segmentation, our time-aware video LLM may struggle with complex/dense temporal reasoning tasks and videos with extreme durations. Our future works will enhance TemporalVLM for tackling complex/dense temporal reasoning tasks and videos with extreme durations.
A.13 Societal Impacts

### v5

@@ Appendix H Effects of Number of Short-Term Clips
Table 12: Effects of number of short-term clips. Best results are in bold, while second best ones are underlined.
Appendix I Results with a More Recent LLM
The performance gain achieved by employing a more advanced LLM as the language decoder in our TemporalVLM model is illustrated in Fig. 9. As LLama 3-8B is pretrained on a significantly larger and diverse dataset, it outperforms Llama 2-7B by a significant margin. The results demonstrate that as LLMs improve, the performance of VLMs, including our TemporalVLM model, will also improve alongside it.
Figure 9: Results with a more recent LLM.
Appendix J Qualitative Results
In addition to Figs. 4 and 5 in the main paper, we provide additional qualitative results in this section. In particular, Figs. 10, 11, and 12 present qualitative results by TemporalVLM in the supervised setting for dense video captioning, video highlight detection, and temporal video grounding respectively. More importantly, qualitative results demonstrating the generalization abilities of TemporalVLM in the zero-shot setting are shown in Fig. 13. Overall, our TemporalVLM model demonstrates promising performance in a variety of temporal reasoning and fine-grained understanding tasks.
@@ Appendix I Results with a More Recent LLM
Appendix K Limitations
Despite promising performance in several temporal reasoning tasks, including dense video captioning, temporal video grounding, video highlight detection, and temporal action segmentation, our time-aware video LLM may struggle with complex temporal reasoning tasks and videos with extreme durations. In addition, our TemporalVLM model often has difficulty dealing with small objects, since it does not include an explicit object detector. Our future works will enhance TemporalVLM for tackling complex temporal reasoning tasks, videos with extreme durations, and dealing with small objects.
Appendix L Societal Impacts